Someone takes 15 minutes to walk up a hill 100m high.His weight is 600N.a)How much work did he do in climbing the hill?

B)what power is developed by his muscles as he climb s?

C)If he climbs up the hill,his muscles can exert 600watts.How long will it take him to getto the top of the hill.


1
Expert's answer
2021-11-26T10:28:30-0500

given quantities:

"t = 15min = 900s"

"h = 100m"

"P = mg = 600N"

"N = 600 \space Watts"

"a) \space A = FS=mgh=600N*100m=60KJ"

"b) \space P = \large\frac{A}{t}" "= 66.6 \space Watts"

"c)" "t = \frac{A}{N}=\frac{60000}{600}=100s = 1min \space 40s"